5.7 quiz
3.) Find an equation of the line that passes through the two given points. Write the equation in slope-intercept form, if possible
x | y
-4| 2
4| 0
8.) There is a linear relationship between a woman's height and the length of her radius bone. It can be stated this way: Height increases by 3.9 inches for each 1-inch increase in the length of the radius. Suppose a 62-inch-tall woman has a 10-inch-long radius bone. Use this information to write a linear equation that relates height h to the length r of the radius. Write the equation in slope-intercept form.
9.) Three months after cleanup began at a dump site, 880 cubic yards of toxic waste had yet to be removed. Two months later, that number had been lowered to 800 cubic yards.
(a) Find an equation that describes the linear relationship between the length of time m (in months) the cleanup crew has been working and the number of cubic yards y of toxic waste remaining. Write the equation in slope-intercept form.
(b) Use your answer to part (a) to predict the number of cubic yards of waste that will still be on the site one year after the cleanup project began.

Answer by stanbon(75887)

Find an equation of the line that passes through the two given points. Write the equation in slope-intercept form, if possible
x | y
-4| 2
4| 0
-------
slope = (2-0)/(-4-4) = 2/-8 = -1/4
-----
Form:
y = mx + b
0 = (-1/4)4 + b
b = 1
y = (-1/4)x + 1
----------------------
8.) There is a linear relationship between a woman's height and the length of her radius bone. It can be stated this way: Height increases by 3.9 inches for each 1-inch increase in the length of the radius. Suppose a 62-inch-tall woman has a 10-inch-long radius bone. Use this information to write a linear equation that relates height h to the length r of the radius. Write the equation in slope-intercept form.
Slope and Point :: m = 3.9 ; (10,62)
y = mx + b
62 = 3.9*10 + b
b = 23
y = 3.9x + 23
------------------
9.) Three months after cleanup began at a dump site, 880 cubic yards of toxic waste had yet to be removed. Two months later, that number had been lowered to 800 cubic yards.
Points:: (3,880) ; (5,800)
(a) Find an equation that describes the linear relationship between the length of time m (in months) the cleanup crew has been working and the number of cubic yards y of toxic waste remaining. Write the equation in slope-intercept form.
slope = (800-880)/(5-3) = -80/2 = -40
y = mx + b
800 = -40*5 + b
b = 1000
y = -40x + 1000
-------------------
(b) Use your answer to part (a) to predict the number of cubic yards of waste that will still be on the site one year after the cleanup project began.
f(12) = -40*12 + 1000 = -480+1000 = 520 cu yds
---------------
Cheers,
Stan H.
---------------